OpenStudy (anonymous):

Evaluate: 5C2 20 25 10 60

OpenStudy (anonymous):

5C2 is: \[\left(\begin{matrix}5 \\ 2\end{matrix}\right) = \frac{5!}{2!(5-2)!} = \frac{5!}{2!3!}\]

OpenStudy (anonymous):

10

OpenStudy (amistre64):

my thought process is as this:\[\] do 5P2; 5*4 = 20\[\] and divide that by 2!, which = 2\[\] 20/2